### v2

# A State-Specific Iterative Decoupling Scheme Based on Perturbation Theory for Low-Energy Electronic States

Abstract
In this work, we introduce a selective and scalable extension of the multi-step Rayleigh–Schrödinger and Brillouin–Wigner (RSBW) perturbative scheme (see Ref. [1]) , designed to efficiently access the low-energy spectrum of molecular systems. The method proceeds by combining successive effective Hamiltonian diagonalizations inspired by second-order Rayleigh–Schrödinger perturbation theory, with a Brillouin–Wigner correction applied to individually optimized states using an updated partitioning of the Hamiltonian. At each step, a zeroth-order state is identified and progressively decoupled from the remaining higher-lying states, thereby enabling a well-conditioned Brillouin–Wigner expansion for the energy correction. In contrast to previous approaches, the method selectively targets a small number of low-lying states, significantly reducing the numerical overhead while maintaining spectroscopic accuracy. The robustness of the method is demonstrated on the LiH and H4 molecules, where accurate excitation energies are obtained for the lowest singlet states using compact model spaces, confirming its potential for realistic applications.
[…]
1 Introduction
Accurately predicting low-lying electronic spectra remains a central challenge in quantum chemistry. Traditional approaches, rooted in the Configuration Interaction (CI) framework, typically rely on diagonalizing the electronic Hamiltonian (or parts of it) to obtain electronic energies and wavefunctions. Without truncation, complete diagonalization yields the so-called Full Configuration Interaction (FCI) expansion, which in principle provides exact solutions within a given basis and thus represents the most accurate approach. However, the exponential scaling of FCI with the system size severely restricts its applicability to small systems, rendering it impractical for realistic molecular studies (although recent works have shown that with extremely large computational resources new records can still be achieved in this direction, see Refs. [2, 3]). To mitigate this limitation, the Complete-Active-Space (CAS) approximation reduces the cost by restricting the expansion to determinants generated within a selected set of active orbitals. This framework underlies widely used approaches such as the CAS Self-Consistent-Field method (and extensions) [4, 5, 6, 7, 8] which have become reference tools of modern computational quantum chemistry. However, the precision of CAS-based methods strongly depends on the choice of an appropriate active space for the calculation [9]. The active space determination is still regarded as an open issue in the electronic structure community and represents the main conceptual limitation of CAS-based methods, even if recent works have focused on developing strategies for automated active space selection[10, 11, 12, 13] and systematic active space expansion [14, 15, 16] (note also the existence of complementary works focusing on quantum information analysis of wavefunctions and orbitals [17, 18, 19, 20]).
As an alternative to brute-force CI expansions, Perturbation Theory (PT) formalisms have proven widely useful for addressing the electronic structure problem. PT can refine an initial calculation into a more accurate treatment, either based on a mean-field reference, leading to the well-known Møller–Plesset approach [21], or on an active-space reference, giving rise to CASPT2 [6, 22] and NEVPT2 [23, 24, 25] methods. PT has also been extensively applied in approaches such as CI using perturbative selection done iteratively, i.e., the so-called CIPSI approach [26, 27, 28, 29], originally proposed to mitigate the computational cost of FCI by guiding selected CI expansions with a perturbative criterion. Building on this idea, the CIPSI scheme was later complemented by the intermediate Hamiltonians approach [30, 31]. These examples illustrate that, beyond serving as a practical tool to refine existing calculations, PT formalisms can also provide a powerful framework for progressively constructing more sophisticated wavefunctions.
Motivated by the flexibility of the PT formalism, in this work we discuss an alternative approach grounded in the so-called Quasi-Degenerate Perturbation Theory (QDPT) [32, 33, 34]. Following QDPT, one traditionally splits the CI space into the model $P$ space (spanned by the $|\psi_{\alpha}\rangle$ configurations with energy $E_{\alpha}$) and the orthogonal $Q$ space (spanned by the $|\psi_{\beta}\rangle$ configurations with energy $E_{\beta}$) and builds the associate effective Hamiltonian operator. Within the Bloch-Rayleigh-Schrodinger flavor of QDPT [32] (which will be central in this work), the resulting form of the second-order effective Hamiltonian within the $P$ space reads as follows:
$\langle\psi_{\alpha}|\hat{H}^{\rm eff}_{2}|\psi_{\alpha^{\prime}}\rangle$ $=\delta_{\alpha\alpha^{\prime}}E_{\alpha^{\prime}}+\langle\psi_{\alpha}|\hat{W}|\psi_{\alpha^{\prime}}\rangle$ (1)
$+\underset{\beta}{\sum}\frac{\langle\psi_{\alpha}|\hat{W}|\psi_{\beta}\rangle\langle\psi_{\beta}|\hat{W}|\psi_{\alpha^{\prime}}\rangle}{E_{\alpha^{\prime}}-E_{\beta}},$
where $\hat{W}$ is the pertubation operator.
Starting from such a space decomposition, recently a two-step scheme has been proposed to combine the canonical Bloch Rayleigh-Schrödinger (RS) perturbation theory (generation of effective Hamiltonians) with a Brillouin-Wigner (BW) correction step. [35, 1] This so-called RSBW method enables the iterative construction of zeroth-order states and their energy corrections, offering a more flexible and scalable alternative to traditional multistate methods. In its previous formulation, the RSBW approach relied on a systematic RS optimization of all reference states, applying a multi-step treatment when needed to refine individual zeroth-order functions. This approach proved effective on model Hamiltonians, revealing that strongly correlated or near-degenerate states may require multiple iterations to achieve optimal decoupling.
However, only a limited number of transitions are of spectroscopic interest in many molecular applications. In such cases, a full optimization of all zeroth-order states, as required in the original RSBW formulation, becomes unnecessarily costly. To address this issue, we propose here a selectiveselective extension of the RSBW method, designed to target only a subset of the low-energy spectrum. By freezing previously optimized states and restricting RS treatments to newly selected ones, this approach significantly reduces the computational overhead while preserving the accuracy and rapid convergence of the BW correction. The state-specific RSBW (SS-RSBW) method provides a more practical route for realistic electronic structure calculations. The remainder of the paper is organized as follows. Section 2 outlines the general procedure underlying the selective RSBW method proposed in this work. In the following sections 3-5, the method is applied to the LiH and H4 molecules, focusing on the lowest singlet states. Those systems serve as a realistic test cases to assess the performance and scalability of the approach.
2 State-specific RSBW scheme
This section details the key steps of the SS-RSBW procedure. The method builds on the foundations of RS and BW perturbation theories, combining iterative effective Hamiltonian constructions with energy corrections tailored to individual electronic states. A key aspect of the approach is the progressive decoupling of selected zeroth-order states from a strategically chosen subset of the remaining zeroth-order states, enabling state-specific and accurate energy predictions while maintaining compact zeroth-order representations. The methodology is introduced through the sequential treatment of the ground and excited states, starting from a reference electronic configurations basis which is gradually refined via a sequence of effective model space optimizations and perturbative corrections.
2.1 Zeroth-order Hamiltonian and initial partitioning
Let $\hat{H}^{(0)}$ denote the Hamiltonian of an arbitrary electronic system, expressed in a finite electronic configurations orthonormal basis $\mathcal{B}^{(0)}=\{|\psi_{k}^{(0)}\rangle\}_{0\leq k\leq N-1}$. In this basis, the Hamiltonian is decomposed as:
$\hat{H}^{(0)}$ $=\sum_{kl}H_{kl}^{(0)}|\psi_{k}^{(0)}\rangle\langle\psi_{l}^{(0)}|$ (2)
$=\underbrace{\sum_{k}H_{kk}^{(0)}|\psi_{k}^{(0)}\rangle\langle\psi_{k}^{(0)}|}_{\hat{H}_{0}^{(0)}}+\underbrace{\sum_{k\neq l}H_{kl}^{(0)}|\psi_{k}^{(0)}\rangle\langle\psi_{l}^{(0)}|}_{\hat{W}^{(0)}}$
$=\sum_{k}E^{(0)}_{k}|\psi_{k}^{(0)}\rangle\langle\psi_{k}^{(0)}|+\sum_{k\neq l}W_{kl}^{(0)}|\psi_{k}^{(0)}\rangle\langle\psi_{l}^{(0)}|.$
In this decomposition, $\hat{H}_{0}^{(0)}$ defines the zeroth-order system, which serves as a reference Hamiltonian, while $\hat{W}^{(0)}=\hat{H}^{(0)}-\hat{H}_{0}^{(0)}$ represents the correlation operator, collecting all off-diagonal contributions. $E_{k}^{(0)}$ is the zeroth-order eigenenergy associated with the eigenstate $|\psi_{k}^{(0)}\rangle$ of $\hat{H}_{0}^{(0)}$. Throughout, we assume the states are energy-ordered, i.e., $E_{k}^{(0)}\leq E_{k+1}^{(0)}$. Our goal is to estimate the energy of the ground state as well as those of the $M$ lowest excited states of the system. The proposed approach proceeds iteratively: we first compute the ground state energy, followed by the first excited state, and so on up to the $M$-th excited state. First, the reference wavefunction is constructed by diagonalizing successive effective Hamiltonians derived from the RS perturbation theory, a step we refer to as the RS treatment. In the second step, this reference state is energy-corrected at second order using a BW perturbative scheme, which captures the residual contributions that are not included in the zeroth-order description. We now detail both steps, starting with the construction of the ground-state reference.
2.2 Ground-state optimization via RS treatments
In our procedure, a zeroth-order eigenstate $|\psi_{k}^{(0)}\rangle$ of the reference Hamiltonian $\hat{H}_{0}^{(0)}$ is deemed an optimal reference state for approximating an eigenstate of the full Hamiltonian $\hat{H}$ if the following condition is fulfilled:
$\rho_{kl}=\left|\frac{\langle\psi_{k}^{(0)}|\hat{W}^{(0)}|\psi_{l}^{(0)}\rangle}{E_{k}^{(0)}-E_{l}^{(0)}}\right|\leq\rho_{\rm min},\quad\forall l\neq k,$ (3)
where $\rho_{\rm min}$ is a predefined threshold.
To construct an optimal zeroth-order state for the ground state, we begin by selecting $|\psi_{0}^{(0)}\rangle$ as the initial reference ($k=0$). This state will be referred to as the Zeroth-Order Candidate State (ZO-CS). All $|\psi_{\alpha}^{(0)}\rangle$ states for which $\rho_{0\alpha}$ (see Eq. 3) exceeds the threshold $\rho_{\rm min}$ are progressively collected to build the model space $P_{0}$. [1] Particular care is required, however, when constructing the second-order effective Hamiltonian defined in Eq. (1). At second order, non-hermitian contributions may arise; to address this, we employ a symmetrization procedure in which the Hamiltonian matrix is averaged with its transpose. Remarkably, the resulting operator can be directly related to the canonical van Vleck QDPT expansion. [36, 33, 37] In addition, some $Q$-space states $|\psi_{\beta}^{(0)}\rangle$ may interact strongly with selected $|\psi_{\alpha}^{(0)}\rangle$ states. To account for this, the model space is further enriched by including the most relevant of these configurations, selected according to a larger threshold $\rho^{\prime}_{\rm min}>\rho_{\rm min}$. Although these additional states are not directly coupled to the ZO-CS, including them in the model space improves the conditioning of the effective Hamiltonian and thus ensures numerical stability during RS optimization.
The diagonalization of the effective Hamiltonian yields a new set of orthonormal eigenstates $\{|\psi_{\alpha}^{\rm RS}\rangle\}$, and an updated basis:
$\mathcal{B}^{(1)}=\{|\psi_{k}^{(1)}\rangle\}=\{|\psi_{\alpha}^{\rm RS}\rangle\}\cup\{|\psi_{\beta}^{(0)}\rangle\},$
where the superscript indicates that a single RS treatment has been performed. The objective of this step is to concentrate part of the electronic effects and to improve the description of the ZO-CS.
The zeroth-order Hamiltonian is then updated as:
$\hat{H}_{0}^{(1)}$ $=\sum_{k}E_{k}^{(1)}|\psi_{k}^{(1)}\rangle\langle\psi_{k}^{(1)}|$ (4)
$=\underbrace{\sum_{\alpha}E_{\alpha}^{\rm RS}|\psi_{\alpha}^{\rm RS}\rangle\langle\psi_{\alpha}^{\rm RS}|}_{\text{updated}}+\underbrace{\sum_{\beta}E_{\beta}^{(0)}|\psi_{\beta}^{(0)}\rangle\langle\psi_{\beta}^{(0)}|}_{\text{unchanged}}.$
The new coupling operator is given by $\hat{W}^{(1)}=\hat{H}^{(1)}-\hat{H}_{0}^{(1)}$, where $\hat{H}^{(1)}$ stands for the representation of the Hamiltonian in the $\mathcal{B}^{(1)}$ basis. Note that the matrix structure of the Hamiltonian remains unchanged within the orthogonal space, whereas the matrix elements involving the model space are updated. Maintaining the energy ordering convention, the new ZO-CS is $|\psi_{0}^{(1)}\rangle$, and the decoupling condition (3) is re-examined to generate an updated model space $P_{0}$. The procedure is repeated until an optimized ZO-CS $|\psi_{0}^{(n_{0}^{\rm RS})}\rangle$ is obtained, where $n_{0}^{\rm RS}$ is the number of required RS steps. This state serves as the zeroth-order reference for the ground-state energy.
2.3 Brillouin–Wigner energy correction
After $n_{0}^{\rm RS}$ RS steps, the zeroth-order ground-state energy is then refined using a second-order Brillouin–Wigner (BW) perturbation correction to reach SS-RSBW energy values. The reference wavefunction is $|\psi_{0}^{(n_{0}^{\rm RS})}\rangle$ and the zeroth-order energy given by $E_{0}^{(n_{0}^{\rm RS})}=\langle\psi_{0}^{(n_{0}^{\rm RS})}|\hat{H}_{0}^{(n_{0}^{\rm RS})}|\psi_{0}^{(n_{0}^{\rm RS})}\rangle$. Assuming that the exact BW energy expansion can be approximated by its second-order value $E_{0}^{\mbox{\tiny SS-RSBW}}$, the BW energy correction reads[32]:
$E_{0}^{\mbox{\tiny SS-RSBW}}$ $=E_{0}^{(n_{0}^{\rm RS})}+\langle\psi_{0}^{(n_{0}^{\rm RS})}|\hat{W}^{(n_{0}^{\rm RS})}|\psi_{0}^{(n_{0}^{\rm RS})}\rangle$
$\quad+\sum_{j\neq 0}\frac{\left|\langle\psi_{0}^{(n_{0}^{\rm RS})}|\hat{W}^{(n_{0}^{\rm RS})}|\psi_{j}^{(n_{0}^{\rm RS})}\rangle\right|^{2}}{E_{0}^{\mbox{\tiny SS-RSBW}}-E_{j}^{(n_{0}^{\rm RS})}}.$ (5)
This equation is either solved iteratively, or by setting $E_{0}^{\mbox{\tiny SS-RSBW}}\approx E_{0}^{(n_{0}^{\rm RS})}$ in the energy denominators.
2.4 Iterative extension to excited states
The same two-step strategy can be sequentially applied to describe excited states. To compute the energy of the first excited state, we start from the optimized zeroth-order basis obtained in the previous step, denoted $\mathcal{B}^{(n_{0}^{\mbox{\tiny RS}})}$. In this basis, the zeroth-order Hamiltonian is rewritten as follows:
$\hat{H}^{(n_{0}^{\mbox{\tiny RS}})}_{0}$ $=\overbrace{E^{(n_{0}^{\mbox{\tiny RS}})}_{0}\,|\Psi^{(n_{0}^{\mbox{\tiny RS}})}_{0}\rangle\langle\Psi^{(n_{0}^{\mbox{\tiny RS}})}_{0}|}^{\text{Optimized}}$
[…]
This partitioning isolates the already optimized part of the spectrum from the remaining states. The former is referred to as the Optimized Zeroth-Order space (OZO-space). At this stage, the OZO-space is restricted to the optimized zeroth-order ground state function $|\psi_{0}^{(n_{0}^{\rm RS})}\rangle$. Let us emphasize that the states within the OZO-space are treated as perturbers in the subsequent RS treatments. That is, they are no longer modified during the optimization of the zeroth-order excited states.
The ZO-CS for targeting the first excited state is the lowest-lying state outside the OZO-space, namely $|\psi_{1}^{(n_{0}^{\rm RS})}\rangle$. Following the previous strategy with identical $\rho_{\rm min}$ and $\rho^{\prime}_{\rm min}$ thresholds, the model space $P_{1}$ is constructed from the $|\psi_{l}^{(n_{0}^{\rm RS})}\rangle$ states with $l\geq 2$. After each RS step, the energies and wavefunctions of the optimized states in the OZO-space are kept frozen, while the remaining ones are energy reordered as needed. The RS optimization is repeated until a new lowest-energy state, orthogonal to the OZO-space, is identified. The resulting optimized state is denoted $|\psi_{1}^{(m_{1}^{\rm RS})}\rangle$, where $m_{1}^{\rm RS}=n_{0}^{\rm RS}+n_{1}^{\rm RS}$ corresponds to the cumulative number of RS steps. This new state is transfered to the OZO-space, and its energy $E^{\mbox{\tiny RSBW}}_{1}$ is computed using the second-order BW correction as defined in Eq. (5). This iterative procedure is applied to construct successively the $M$ lowest eigenstates of the full Hamiltonian. The newly $i$-th zeroth-order state computed after $n_{i}^{\rm RS}$ RS steps is added to the OZO space. This sequential decoupling approach ensures that each targeted state is treated independently, while maintaining the structure of OZO-space states. However, it should be noted that the algorithm does not guarantee that the order in which states are optimized may differ from the exact one. For instance, the second excited state may occasionally be obtained before the first one. The final set of SS-RSBW energies is systematically reordered in ascending energy at the end of the calculation. A compact summary of the full sequence of operations defining the SS-RSBW scheme used in this work is presented in Algorithm 1.
Algorithm 1 $S$-RSBW procedure for computing the energies of $M$ targeted states. In practice, the procedure allows to generate the lowest states starting from the Hartree-Fock $\hat{H}^{(0)}$ Hamiltonian.
1:Step 1: Initialization
2:Set thresholds: $\rho_{\rm min}\leftarrow\text{e.g., }0.1$,  $\rho^{\prime}_{\rm min}\leftarrow\text{e.g., }0.5$
3:Define $\hat{H}_{0}^{(0)}$ such that $(H_{0}^{(0)})_{ij}=\delta_{ij}H_{ij}^{(0)}$
4:Set $\hat{W}^{(0)}\leftarrow\hat{H}^{(0)}-\hat{H}_{0}^{(0)}$
[…]
6:Set $i\leftarrow 0$ $\triangleright$ Index of the targeted state
7:Set $m_{i}^{\rm RS}\leftarrow 0$ $\triangleright$ Cumulative number of RS iterations for state $i$
8:loop over $i=0,\dots,M-1$
9:  Step 2: Iterative RS treatment
10:  Build model space $P_{i}=\left\{|\psi_{\alpha}^{(m_{i}^{\rm RS})}\rangle\right\}$
[…]
12:   Build orthogonal $Q_{i}$-space
13:   Build and symmetrize $\hat{H}_{\rm eff}^{(2)}$
14:   $\left\{E_{\alpha}\right\}\leftarrow$ Diagonalize $\hat{H}_{\rm eff}^{(2)}$
15:   $m_{i}^{\rm RS}\leftarrow m_{i}^{\rm RS}+1$
16:   Update $\mathcal{B}^{(m_{i}^{\rm RS})}$, $\hat{H}^{(m_{i}^{\rm RS})}$
17:   Update $\langle\psi_{\alpha}^{(m_{i}^{\rm RS})}|\hat{H}_{0}^{(m_{i}^{\rm RS})}|\psi_{\alpha}^{(m_{i}^{\rm RS})}\rangle\leftarrow\left\{E_{\alpha}\right\}$
18:   Set $\hat{W}^{(m_{i}^{\rm RS})}\leftarrow\hat{H}^{(m_{i}^{\rm RS})}-\hat{H}_{0}^{(m_{i}^{\rm RS})}$
19:   Sort basis states in ascending energy
20:   Build $P_{i}$ in the new basis $\mathcal{B}^{(m_{i}^{\rm RS})}$
21:  end while
22:  Step 3: BW correction
23:  Compute $E^{\mbox{\tiny RSBW}}_{i}$ using Eq. (7)
24:  Add $|\psi_{i}^{(m_{i}^{\mbox{\tiny RS}})}\rangle$ to the OZO-space
25:  Set $i\leftarrow i+1$, $m_{i}^{\rm RS}\leftarrow m_{i-1}^{\rm RS}$
26:end loop
27:Sort SS-RSBW energies in ascending energy order
3 Numerical tools
We numerically tested our approach on two prototypical molecular systems, LiH and H4 ring, using STO-3G atomic orbitals (AOs): specifically, the $1s$ orbital for hydrogen and the $1s$, $2s$, and $2p$ orbitals for lithium. The latter are used to generate canonical molecular orbitals (MOs) via Restricted Hartree-Fock (RHF) calculations implemented in the Psi4 package. [38] Then, the QuantNBody package [39, 40] allows to transform the one- and two-electron integrals from the AOs to the MOs basis, generate the Slater determinants basis set, and construct the electronic Hamiltonian matrix. To isolate the singlet subspace, a configuration state functions (CSFs) basis is constructed from spin-adapted linear combinations of Slater determinants. This is done by building and diagonalizing the $\hat{S}^{2}$ operator over the determinants space within QuantNBody, yielding an orthonormal spin-adapted basis. The Hamiltonian is then transformed into this basis, becoming block-diagonal, and the singlet CSFs define the zeroth-order basis $\mathcal{B}^{(0)}$ used throughout the procedure.
4 Application to lithium hydride
The proposed SS-RSBW method was first applied to compute the low-lying singlet electronic states of the lithium hydride (LiH) molecule. Specifically, we focused on the ground and the first two excited singlet states, as a function of the internuclear distance $R$. The FCI space is spanned by the Slater determinants arising from distributing four electrons in twelve spin-orbitals, leading to $N=\binom{12}{4}=495$ determinants. The FCI ground-state energy profile exhibits a minimum at $R_{\rm eq}=1.55\penalty 10000\ \AA$ , showing a small deviation from the reported equilibrium bond distance of $1.60\penalty 10000\ \AA$ . Despite this discrepancy, the STO-3G basis set was retained to generate the reference FCI calculations. To investigate the evolution of the electronic structure across both compressed and stretched bonding regimes, the internuclear distance $R$ is varied from $0.75$ to $2.20\penalty 10000\ \AA$ . The zeroth-order basis $\mathcal{B}^{(0)}$ is defined by the $N=105$ singlet CSFs.
4.1 ZO-CSs optimization: RS treatments
The optimization of the ZO-CSs was carried out using the RS scheme introduced in Section 2.2. The selection of states included in the model space is governed by the two thresholds $\rho_{\rm min}=0.1$ and $\rho^{\prime}_{\rm min}=0.5$, used across all targeted states.
Ground state: The initial ZO-CS for the ground state was selected as the lowest-lying CSF. The dimensions of the model spaces during the successive RS treatments used to optimize this reference state are shown in Figure 1.
[…]
For internuclear distances $R<2.0\penalty 10000\ \AA$ , the model space dimension is less than two, and the number of RS steps is $m_{0}^{\rm RS}=1$. As expected, the dimension of the model space as well as $m_{0}^{\rm RS}$ increase in the bond-stretched regime $R>2.0\penalty 10000\ \AA$ where static correlation effects must be included. To assess the impact of RS optimization, the decompositions of the exact ground-state FCI wavefunction are compared for the initial and optimized zeroth-order bases in Figure 2.
[…]
The decomposition of the ground-state wavefunction in the initial CSF basis (Figure 2-a) exhibits a strong dominance of the ZO-CS which is primarily composed of the RHF determinant. This observation is consistent with the well-established accuracy of the RHF description for ground-state properties in weakly correlated regimes [41]. However, as the LiH bond length $R$ increases, the wavefunction acquires a pronounced multiconfigurational character (i.e., static correlation), as evidenced by the growing number of significant contributions from other CSFs in the $\mathcal{B}^{(0)}$ basis (Figure 2-a(III)). The picture is significantly modified in the $\mathcal{B}^{(n_{0}^{\mbox{\tiny RS}})}$ basis. As seen in Figure 2-b, the projection onto the optimized ZO-CS dominates as soon as the $\mathcal{B}^{(n_{0}^{\mbox{\tiny RS}})}$ basis is constructed. The sequence of RS treatments in this region systematically incorporates the most relevant configurational components needed for an accurate zeroth-order state that offers a significantly improved description of the ground state compared to the initial CSF-based ZO-CS.
First excited singlet state: A similar analysis was carried out for the first excited singlet state. The initial ZO-CS is selected from the updated basis $\mathcal{B}^{(n_{0}^{\rm RS})}$ as the lowest-energy state orthogonal to the OZO-space, namely $|\psi_{1}^{(n_{0}^{\rm RS})}\rangle$. The associated model space dimensions $P_{1}$ and wavefunction decompositions are presented in Figures 3 and 4, respectively.
[…]
Figure 4: Decomposition of the exact first excited singlet state on (a) the initial CSF basis $\mathcal{B}^{(0)}$, and (b) the optimized basis $\mathcal{B}^{(m_{1}^{\rm RS})}$. In the optimized basis, the projection onto the candidate state is indexed by $k=0$, while the projections onto the other states are ordered in decreasing absolute value. In the CSFs basis, projections are ordered by decreasing absolute value. For further details, refer to the caption of Fig. 2.
First, the model space dimensions $P_{1}$ involved in the successive RS treatments remain rather small whatever the internuclear distance (see Figure 3). Interestingly, the starting ZO-CS $|\psi_{1}^{(n_{0}^{\rm RS})}\rangle$, benefits from the refinements performed through the optimization of the ground state. Despite its higher multiconfigurational character in the $\mathcal{B}^{(0)}$ basis set (see Figure 4, panel (a)), the computational effort required for its optimization is significantly reduced when the $\mathcal{B}^{(n_{0}^{\rm RS})}$ basis set is used. Finally, Figure 4-b shows that the first excited singlet state exhibits a clear dominant projection on the ZO-CS, illustrating the ability of the RS treatment to effectively concentrate the essential electronic information into the selected reference state.
Second excited singlet state: The ZO-CS for the second excited singlet state was optimized using the same sequential RS strategy, with $|\psi_{2}^{(m_{1}^{\rm RS})}\rangle$ ($m_{1}^{\rm RS}=n_{0}^{\rm RS}+n_{1}^{\rm RS}$) selected as the initial reference state. As highlighted for the first excited state, the preceding RS treatments may contribute to a pre-conditioning of the state $|\psi_{2}^{(m_{1}^{\rm RS})}\rangle$, thereby reducing the computational effort required for its optimization. This is reflected in the dimensions of the model spaces $P_{2}$ which remain relatively small, as illustrated in Figure 5. The observed behavior for the optimization of the ZO-CS for the second excited singlet state closely parallels that of the first excited state: the RS treatments reliably yield a dominant component in the optimized state, even in strong correlation regimes.
Figure 5: Model spaces dimensions $P_{2}$ used throughout the optimization of the ZO-CS for the second excited singlet state. For further details, refer to the caption of Figure 1.
These results underscore the efficiency of the RS procedure in capturing multiconfigurational effects within a single optimized vector, thereby enabling a compact yet accurate representation of both ground and excited electronic states across a broad range of correlation regimes.
Figure 6: Decomposition of the exact first excited singlet state on (a) the initial CSFs basis $\mathcal{B}^{(0)}$, and (b) the optimized basis $\mathcal{B}^{(m_{2}^{\rm RS})}$, where $m_{2}^{\rm RS}$ is the total number of RS steps. In the optimized basis, the projection onto the candidate state is indexed by $k=0$, while the projections onto the other states are ordered in decreasing absolute value. In the CSF basis, projections are ordered by decreasing absolute value. For further details, refer to the caption of Figure 2.
4.2 Energy evaluation via BW correction
The SS-RSBW energies of the singlet states were computed by applying second-order BW corrections to the zeroth-order energies of the optimized states $\{|\psi_{k}^{(m_{k}^{\rm RS})}\rangle\}$. The corrected energy $E_{k}^{\mbox{\tiny SS-RSBW}}$ is defined by:
$E_{k}^{\mbox{\tiny SS-RSBW}}$ $=E_{k}^{(m_{k}^{\rm RS})}+\langle\psi_{k}^{(m_{k}^{\rm RS})}|\hat{W}^{(m_{k}^{\rm RS})}|\psi_{k}^{(m_{k}^{\rm RS})}\rangle$
$\quad+\sum_{j\neq k}\frac{\left|\langle\psi_{k}^{(m_{k}^{\rm RS})}|\hat{W}^{(m_{k}^{\rm RS})}|\psi_{j}^{(m_{k}^{\rm RS})}\rangle\right|^{2}}{E_{k}^{\mbox{\tiny SS-RSBW}}-E_{j}^{(m_{k}^{\rm RS})}}.$ (7)
where $E_{k}^{(m_{k}^{\rm RS})}=\langle\psi_{k}^{(m_{k}^{\rm RS})}|\hat{H}_{0}^{(m_{k}^{\rm RS})}|\psi_{k}^{(m_{k}^{\rm RS})}\rangle$.
Figure 7: SS-RSBW energies $\{E^{\text{SS-RSBW}}_{0\text{–}2}\}$ of the three lowest singlet states of LiH as a function of the internuclear distance $R$. Exact FCI results $\{E^{\mathrm{exact}}_{0\text{–}2}\}$ are shown for comparison.
The quality of the computed energy levels obtained from both RS and SS-RSBW methods, was assessed using the root-mean-square (RMS) error, defined as
$\Delta^{\text{method}}=\sqrt{\frac{1}{N_{R}}\sum_{i_{R}=1}^{N_{R}}\left(E^{\text{method}}_{i_{R}}-E^{\rm FCI}_{i_{R}}\right)^{2}},$
where the upperscript "method" is either RS or SS-RSBW, and $N_{R}=51$ denotes the number of grid points used to sample the internuclear distance $R$. The RMS errors $\Delta^{\rm RS}$ and $\Delta^{\text{SS-RSBW}}$ for the three computed singlet energy levels are reported in Table 1.
Energy Level $E_{0}$ $E_{1}$ $E_{2}$
$\Delta^{\rm RS}$ (mHa) $1.0$ $2.3$ $0.1$
$\Delta^{\text{SS-RSBW}}$ (mHa) $0.07$ $0.08$ $0.003$
Table 1: Root-mean-square errors (mHa) for the three lowest singlet energy levels of LiH. The first row reports errors from RS treatments, while the second row corresponds to the full SS-RSBW results. Reference values are obtained from FCI calculations. All computations are performed using the STO-3G AO basis set.
Table 1 suggests that the RS treatments already provide remarkably accurate zeroth-order energies over the bond-length range, reaching spectroscopic accuracy (RMS $\sim$ 1 mHa). This result underscores the robustness of the underlying partitioning and confirms that the reference states are both accurate and systematically improvable. Building on this, the full SS-RSBW scheme further reduces the errors: for all three singlet states, the RMS deviations from the exact FCI energies remain below 0.1 mHa. The algorithm converges rapidly, requiring less than five iterations to solve Eq. (7) for all geometries. This highlights both the quality of the RS-optimized zeroth-order states and the overall efficiency of the scheme for describing low-lying singlet states in correlated electronic systems. Besides, the enhanced convergence suggests that higher-order BW energy expansion might be included, shifting away the size-consistency drawback of the BW theory.
5 Application to the H4 ring
Figure 8: Geometry of the ring H4 model with angular parametrization. The radius $R$ is fixed at 1 Å, and the angle $\theta$ varies from $60^{\circ}$ to $120^{\circ}$.
This section addresses the application of the SS-RSBW method to the H4 ring model. The system consists of four hydrogen atoms arranged on a circle of fixed radius R=1.0ÅR=1.0\ $\text{\AA }$, while keeping a C2 axis with a characteristic $\theta\in[60^{\circ},120^{\circ}]$ angle (see Figure 8). This controlled deformation of the system provides a simple yet stringent testbed to probe the emergence and removal of orbital degeneracies associated with the transition between the $D_{2h}$ and $D_{4h}$ point groups. Such geometries induce strong static correlation effects that are poorly described by single-reference methods, making H4 a classical benchmark for assessing the robustness of multireference approaches [42, 43, 44, 45]. The FCI space built on four electrons in eight spin-orbitals comprises $\binom{8}{4}=70$ Slater determinants combined into 20 spin-adapted singlet CSFs.
5.1 ZO-CS optimization: RS treatments
The optimization of ZO-CSs for ground and excited states through RS treatments follows the procedure presented earlier, using the thresholds values $\rho_{\min}=0.05$ and $\rho^{\prime}_{\min}=0.2$.
Ground state:
Figure 9: Optimization of the ZO-CS for the ground singlet state: (a) dimensions of the model spaces $P_{0}$ as a function of the angle $\theta$; (b) and (c) decompositions of the FCI ground state over the initial CSF basis $\mathcal{B}^{(0)}$, and the optimized zeroth-order basis $\mathcal{B}^{(n_{0}^{\rm RS})}$, respectively, shown for (I) rectangular and (II) square geometries. The projection onto the ZO-CS is indexed by $k=0$, and others are ordered by descending absolute values.
As seen in Figure 9-a, a single iteration ($n_{0}^{\rm RS}=1$) is necessary to optimize the ZO-CS whatever the $\theta$ value. The strong multi-configurational character is reflected by the dimension of the model space that remains relatively large as compared to the dimension of the full CSFs basis set. Then, special attention was dedicated to the rectangular $\theta=60^{\circ}$ (I) and square $\theta=90^{\circ}$ (II) geometries. As seen in Figure 9-b, the decompositions of the ground state over $\mathcal{B}^{(0)}$ reveal significant contributions arising from several CSFs, with a more pronounced static correlation manifestation for the square geometry (see Figure 9-b(II)). The $D_{4h}$ symmetry induces orbital degeneracies which, in the singlet CSF basis, manifest as a broader expansion of the wavefunction. Finally, Figure 9-c highlights a dominant projection of the FCI ground state onto the optimized ZO-CS for both geometries, thus confirming the reliability of the constructed reference state with the RS treatment.
Figure 10: Optimization of the ZO-CS for the first excited singlet state: (a) dimensions of the model spaces $P_{1}$ as a function of $\theta$; (b) and (c) decomposition of the FCI first excited state over the initial CSF basis and the optimized zeroth-order basis $\mathcal{B}^{(m_{1}^{\rm RS})}$, respectively, shown for (I) rectangular and (II) square geometries. In the optimized basis, the projection onto the candidate state is indexed by $k=0$, while the projections onto the other states are ordered in decreasing absolute value. In the CSF basis, projections onto all states are ordered by decreasing absolute value.
Excited singlet states: For most geometries, the ZO-CS extracted from the previously optimized basis $\mathcal{B}^{(n_{0}^{\rm RS})}$ already provides an optimal description, making additional RS treatments unnecessary and reducing the dimensions of the model space $P_{1}$ as compared to $P_{0}$ (see Figure 10-a). However, in the vicinity of $\theta=66^{\circ}$ (and its symmetric counterpart $\theta=114^{\circ}$), the dimension of the model space increases with two (quasi-)degenerated excited singlet states. This point will be addressed later in this section. As for the ground state, the decompositions of the first excited state on $\mathcal{B}^{(0)}$ shown in Figure 10-b extend over several components, particularly in the square geometry (see Figure 10-b(II)). Finally, Figure 10-c stresses the compact expansion of the FCI excited state onto the optimized ZO-CS, favouring the subsequent BW expansion.
Similar conclusions can be drawn for the second excited state. First, the model space is significantly reduced as compared to the decomposition onto the $\mathcal{B}^{(0)}$ basis. Second, the optimized ZO-CS strongly overlap with the corresponding FCI eigenstate. Therefore, the effectiveness of the RS procedure in constructing a representative reference state is confirmed for the second singlet excited state as well.
5.2 BW-corrected energies
Figure 11: SS-RSBW energies $\{E^{\text{SS-RSBW}}_{0\text{–}2}\}$ of the ground and two lowest singlet excited states of the H4 ring as a function of $\theta$. Exact FCI results $\{E^{\mathrm{exact}}_{0\text{–}2}\}$ are shown for comparison. The first and second excited states crossings for $\theta=66^{\circ}$ (and symmetric counterpart $\theta=114^{\circ}$) are accurately reproduced by the SS-RSBW procedure.
Energy Level $E_{0}$ $E_{1}$ $E_{2}$
$\Delta^{\rm RS}$ (mHa) $0.4$ $0.02$ $0.08$
$\Delta^{\text{SS-RSBW}}$ (mHa) $0.04$ $0.01$ $0.01$
Table 2: Root-mean-square errors (mHa) for the three lowest singlet energy levels of H4. The first row reports errors from RS treatments, while the second row corresponds to the SS-RSBW results. References are FCI values.
Figure 11 shows the FCI energies of the ground state and the two lowest singlet excited states as a function of $\theta$, together with the SS-RSBW results (dashed line). The SS-RSBW energies were obtained from a second-order BW expansion and converged with less than five iterations (see Eq. 5). Overall, the agreement is excellent (see Table 2) for both SS-RSBW and RS energies across the entire angular range considered.
For $\theta=66^{\circ}$ and $\theta=114^{\circ}$, the two excited states are degenerate, a feature which is reproduced with high accuracy by our approach. In practice, the procedure triggers a local enrichment of the model space during the optimization of the ZO-CS for the first excited state (see Figure 10-a), thereby explicitly including the CSFs responsible for the coupling of the two states. This targeted enlargement of the model space improves the conditioning of the BW correction and enables the method to recover two numerically degenerate levels, whereas a naïve fixed-reference perturbative strategy would likely fail. This behavior has also been analyzed in previous works on this method [1].
6 Conclusion
We have introduced and applied a perturbative framework that combines an iterative Rayleigh–Schrödinger (RS) optimization of zeroth-order states with second-order Brillouin–Wigner (BW) correction. This two-step SS-RSBW method enables the construction of compact and refined reference wavefunctions that concentrate the multiconfigurational character of both ground and excited states. By targeting individual eigenstates successively, and optimizing one at a time their zeroth-order representation, the method offers a controlled and accurate alternative to traditional multi-state approaches.
Applications to the LiH and H4 ring molecules demonstrate the effectiveness of the proposed scheme in both accuracy and computational efficiency. Throughout the geometry changes, the optimized zeroth-order candidate states consistently exhibit dominant overlap with the exact FCI solutions. Whatever the geometries, the BW-corrected energies for the three lowest singlet states of LiH and H4 show very good agreement with FCI reference values, with root-mean-square (RMS) errors consistently lower than $10^{-4}$ Hartree. These results underscore the ability of the RS treatment to isolate relevant part of correlation effects into low-dimensional model spaces, enabling efficient, yet highly accurate, energy predictions even in strongly correlated regimes. While targeted molecules remain relatively small systems for which full FCI calculations are still tractable, it is to be considered as a valuable benchmark for demonstrating the reliability and underlying principles of the SS-RSBW strategy. The compactness of the model spaces and the rapid convergence observed in this study suggest that the method can scale efficiently with system size and complexity. Future work will explore the extension of this approach to larger molecular systems and more challenging electronic structures, including multi-state potential energy surfaces relevant to photophysical and photochemical processes.
7 Acknowledgments
This work was supported by the ANR (PRC ANR-2023-CE07-0035 AtropoPhotoCat). The authors wish to thank Prof. Celestino Angeli and Dr. Jean-Paul Malrieu for helpful discussions regarding the transferability of the method.